# Learn2Reg 2024: New Benchmark Datasets Driving Progress on New Challenges

### v2

@@ Abstract
Medical image registration is critical for clinical applications, and fair benchmarking of different methods is essential for monitoring ongoing progress. To date, the Learn2Reg 2020-2023 challenges have released several complementary datasets and established metrics for evaluations. However, these editions did not capture all aspects of the registration problem, particularly in terms of modality diversity and task complexity. To address these limitations, the 2024 edition introduces three new tasks, including large-scale multi-modal registration and unsupervised inter-subject brain registration, as well as the first microscopy-focused benchmark within Learn2Reg. The new datasets also inspired new method developments, including invertibility constraints, pyramid features, keypoints alignment and instance optimisation.
doi:
10.59275/j.melba.2024-AAAA
@@ 2.1 Medical Image Registration Challenges
Several notable medical image registration challenges have significantly advanced the field. The EMPIRE10 challenge focused on lung CT registration, offering comprehensive evaluations including landmark distances and Jacobian determinants (Murphy et al., 2011). The CuRIOUS challenge addressed intra-operative ultrasound and MRI registration (Xiao et al., 2019), while ANHIR specialized in histological image registration (Borovec et al., 2020), each attracting multiple participating teams and employing diverse evaluation metrics. The Continuous Registration Challenge, co-hosted with WBIR 2018, aimed to unify multiple registration tasks but was limited by its integration requirements, specifically the mandatory use of the SuperElastix C++ API framework (Marstal et al., 2019). More recently, the ACROBAT challenges in 2022 and 2023 addressed whole-slide image registration for breast cancer tissue, utilizing large datasets and providing valuable insights into algorithm performance (Weitz et al., 2024). These challenges have collectively advanced the development and benchmarking of medical image registration algorithms, each contributing unique datasets and frameworks that have influenced Learn2Reg tasks and evaluation.
@@ 2.2 Learn2Reg to Date
Since its inception, the Learn2Reg challenge has served as a key platform for advancing medical image registration through open benchmarking and community collaboration. Over multiple editions, the challenge has brought together researchers worldwide to tackle diverse registration problems across various imaging modalities and anatomical sites. This section provides a brief overview of the challenge’s objectives, datasets, and key developments up to the present edition. The Learn2Reg series began with the Learn2Reg Tutorial at MICCAI 2019 in Shenzhen, which introduced the community to emerging deep learning approaches for medical image registration through expert lectures and practical sessions (Learn2Reg Organisers, 2019). This foundational event set the stage for the challenge’s collaborative and multidisciplinary focus.
The 2020 and 2021 editions of Learn2Reg were the initial official challenges, introducing six datasets covering multiple anatomies—including brain, abdomen, and thorax—and imaging modalities such as MRI, CT, and ultrasound. Both editions were held as community challenges with workshops at the respective MICCAI conferences. They followed a Type 1 challenge format, providing training images and labels, with test images available but test labels withheld for evaluation. These editions also introduced the challenge’s open-source evaluation and ranking code, as well as baseline methods, enabling standardized and reproducible benchmarking. Together, they attracted over 20 international teams submitting more than 65 methods, offering comprehensive insights into the performance of conventional and deep-learning registration approaches across multiple clinically relevant tasks (Hering et al., 2022).
In 2022 and 2023, Learn2Reg held two further challenge workshops at MICCAI, focusing on the continuation of the initial benchmarks while expanding the scope of the challenge. These editions introduced a new dataset, NLST (CT Lung Screening Registration Data), and pushed towards Type 3 challenges—where both training and test images, along with their labels, remain hidden during evaluation to better simulate real-world conditions.
In parallel, the OncoReg spin-off challenge was launched at MICCAI 2023, followed by a standalone workshop in early 2024. It introduced the ThoraxCBCT dataset, which includes longitudinal lung CBCT and FBCT scans collected from real clinical settings. OncoReg specifically targets oncological applications, aiming to advance AI-driven medical image registration to monitor lung cancer progression and improve radiotherapy planning. Through multidisciplinary collaboration, including expert anatomical annotations and advanced AI methods, OncoReg addresses the pressing need for robust registration tools in lung cancer research and treatment (Heyer et al., 2025). Overall, the 2022–2023 editions further consolidated Learn2Reg’s role as a platform for diverse, clinically relevant datasets and moved closer to real-world deployment of medical image registration solutions. The evolution from foundational benchmarks to clinically focused and modality-diverse challenges reflects Learn2Reg’s growing impact in the field. The 2024 edition continues this trajectory, expanding into new imaging domains and addressing increasingly complex registration problems.
@@ 3.1 Challenge Organisation
The Learn2Reg 2024 challenge was organised by a dedicated core team led by Mattias Heinrich (University of Lübeck), who coordinated all aspects of the event, including website maintenance, evaluation infrastructure, and the associated workshop. The challenge consisted of three tasks, each developed and maintained by a separate expert team: ReMIND2Reg (Harvard Medical School, led by Reuben Dorent), LUMIR (Johns Hopkins University, led by Junyu Chen), and COMULISglobe SHG/BF (COMULISglobe community, led by Andreas Walter, Hochschule Aalen, (Walter et al., 2021; Rudraiah et al., 2024; Walter et al., 2020)). These tasks addressed diverse challenges in biomedical image registration and are described in detail in later sections.
As in previous editions, Learn2Reg 2024 followed a two-phase structure hosted on the Grand Challenge platform [^1] https://learn2reg.grand-challenge.org/. During the validation phase, participants were given access to training and validation data to develop and test their methods locally. In the subsequent test phase, participants submitted their final algorithms, which were evaluated on hidden test datasets. The evaluation was performed using a standardized, open-source framework [^2] https://github.com/MDL-UzL/L2R, ensuring fairness, reproducibility, and transparency. Performance was assessed using task-specific quantitative metrics, with results published on public leaderboards. The ranking scheme is based on the ranking scheme of the Medical Decathlon (statistical comparisons of method performance across metrics, aggregated into overall task scores) (Antonelli et al., 2022).
@@ 3.2.1 ReMIND2Reg
The challenge task is to register post-resection 3D iUS volumes to either contrast-enhanced T1-weighted (ceT1) or T2-weighted MRI, depending on availability for each case. Submitted methods must be able to handle missing modalities and account for large deformations, resections, and noise inherent in iUS data. The dataset is derived from the ReMIND collection (Juvekar et al., 2023) at Brigham and Women’s Hospital. Specifically, the ReMIND2Reg dataset includes 99 training cases with 99 3D iUS, 93 ceT1, and 62 T2 scans, 4 validation cases with 4 3D US, 4 ceT1, and 4 T2, and 10 test cases with 10 3D US, 10 ceT1 and 10 T2.
@@
Full data and acquisition protocols are described in (Juvekar et al., 2023). The dataset is shared under a CC BY 4.0 license and is available via TCIA [^3] https://doi.org/10.1101/2023.09.14.23295596.
@@ 3.2.2 LUMIR
Participants are required to submit deformation fields aligning specified image pairs. The performance is assessed via Dice similarity and Hausdorff-95 distance on segmentation labels, target registration error on anatomical landmarks, and the fraction of non-diffeomorphic volumes (NDV) (Liu et al., 2024b) to measure deformation plausibility. Baseline code and evaluation scripts are provided online [^4] https://github.com/JHU-MedImage-Reg/LUMIR_L2R. An introductory video is also available [^5] https://cloud.imi.uni-luebeck.de/s/tm9DEFiDa9XH35k/download/Learn2Reg_LUMIR_JChen.mp4.
This paper provides a general overview of the LUMIR task in the context of Learn2Reg 2024. A more in-depth analysis of the LUMIR challenge design, results, and implications for developing foundational registration models can be found in (Chen et al., 2025). Relevant datasets and prior methods are further detailed in (Dufumier et al., 2022; Taha et al., 2023; Marcus et al., 2007; Liu et al., 2024b).
@@ 3.2.3 COMULISglobe SHG/BF
Further details on the underlying imaging and dataset are available in (Keikhosravi et al., 2020). More information on the goals of the COMULISglobe initiative and the broader community can be found at the COMULISglobe website [^6] https://www.comulis.eu/.
@@ 3.2.4 NLST
The challenge consists of two phases. In Phase 1, participants developed and tested algorithms locally using 110 training and 22 validation cases, all sourced from the National Lung Screening Trial (NLST) (Team, 2011)  [^7] https://wiki.cancerimagingarchive.net/display/NLST/National+Lung+Screening+Trial. Each case consists of a pair of thoracic CT scans acquired either longitudinally (screening cohort) or across different respiratory phases (radiotherapy cohort). Automatic keypoint correspondences (Heinrich et al., 2015) and lung masks are provided to support model development. In Phase 2, the top-performing teams were invited to submit training Docker containers for remote execution on a hidden, extended dataset comprising 1300 training, over 100 validation, and more than 200 test cases. This second part of the dataset includes additional annotations not publicly released.
@@ 4.1.1 ReMIND2Reg
Figure 3: Results for the ReMIND2Reg task. Target Registration Error (TRE) on manual landmarks and TRE30 (the TRE computed on the 30% of landmarks with the largest initial errors) are reported for all methods.
@@
During the validation phase of the ReMIND2Reg task, a total of thirteen teams submitted results, along with two baseline methods provided by the challenge organizers. The task’s complexity was reflected in the exceptionally high number of validation entries, with 253 submissions indicating that teams required extensive iterative optimization to handle the challenging multimodal ultrasound-to-MRI registration problem. In the test phase, participation reduced to five teams plus one baseline method. This reduction highlights the technical difficulty of registering post-resection intraoperative ultrasound to preoperative MRI, particularly given the presence of brain shift and tissue deformation. One baseline was provided for the ReMIND2Reg test phase. The submitting teams included junyi-wang, VROC, next-gen-nn, IGRTS, and SamPLe.
@@
niftyreg, which employed the NiftyReg open-source software package [^8] https://github.com/KCL-BMEIS/niftyreg/wiki implementing a symmetric block-matching registration approach. This method used iterative block-matching to establish point correspondences between ultrasound and MRI images, followed by least trimmed squares (LTS) regression to determine transformation parameters. To identify informative regions, the algorithm divided images into uniform blocks and selected those with the highest intensity variance. Block similarity was assessed using normalized cross-correlation, and matching was performed based on these scores. To ensure inverse consistency, the method averaged directional transformation matrices at each iteration, maintaining symmetry throughout the registration process (Drobny et al., 2018).
@@
junyi-wang developed a coarse-to-fine pipeline combining imaging style transfer with traditional registration techniques. Their approach used 3D CycleGAN to generate synthetic T1-style images from ultrasound data, enabling unified signal distributions across modalities. The registration process employed hierarchical block matching via NiftyReg for affine transformation estimation, followed by a pretrained SynthMorph model for local deformation inference around resection areas. VROC employed its fully differentiable variational registration framework (Werner et al., 2014; Ehrhardt et al., 2015) [^9] https://github.com/IPMI-ICNS-UKE/vroc, combining task-specific preprocessing with rigid registration optimization. For the ReMIND2Reg task, VROC applied Gaussian smoothing and threshold-based clipping of ultrasound intensity values during preprocessing. The registration pipeline performed initial vector field estimation via global normalized cross-correlation (NCC) loss optimization, followed by fine-tuning using normalized gradient fields (NGF) loss optimization to handle the challenging multimodal ultrasound-to-MRI registration scenario.
@@
next-gen-nn addressed the ReMIND2Reg challenge using an unsupervised multimodal registration method based on Multilevel Correlation Balanced Optimization (MCBO), building upon the ConvexAdam framework (Siebert et al., 2024). The approach was specifically designed to handle two key difficulties: the absence of supervised labels and the presence of large deformations between pre-operative MRI and intra-operative ultrasound images. For paired multimodal moving images (ceT1 and T2), the method adopted maximum fusion strategies to optimize registration performance across the challenging modality gap. IGRTS developed a topological higher-order Markov Random Field approach for deformable registration. Their method utilized multiscale multimodal Gabor attribute vectors to characterize MRI and ultrasound images, with the 3D deformation field’s topological framework constructed through Jacobian determinants derived from higher-order cliques. The optimization employed a multi-resolution QPBO (Quadratic Pseudo-Boolean Optimization) strategy to handle the complex higher-order MRF energy function.
@@ 4.1.2 LUMIR
Figure 4: Results for the LUMIR task. The table presents final test set results for all participating teams and baseline methods. Reported metrics include Target Registration Error (TRE) on manual landmarks, Dice-Sørensen coefficient (DSC) and 95th percentile Hausdorff distance (HD95) on brain label maps, as well as the non-diffeomorphic volume (NDV) of displacement fields. The accompanying box plot provides a visual overview of DSC performance distribution across all methods.
@@
A distinctive aspect of the LUMIR challenge was its extensive post-challenge analysis, documented in detail in ”Beyond the LUMIR challenge: The pathway to foundational registration models” (Chen et al., 2025). This companion paper provides comprehensive validation with particular emphasis on foundational registration models, zero-shot evaluation, and out-of-domain generalization tasks. The present section summarizes the key findings and participating methods, while readers are referred to the companion paper for in-depth technical analysis and broader implications for developing generalizable brain MRI registration frameworks. The LUMIR challenge included eight baseline methods spanning classical optimization-based approaches and modern deep learning frameworks. ANTsSyN employed symmetric, time-varying diffeomorphic registration using cross-correlation within the space of diffeomorphic maps (Avants et al., 2008). deedsBCV utilized discrete optimization with Markov random field strategies and minimum spanning trees to handle sliding motions efficiently (Heinrich et al., 2012). SynthMorph implemented a contrast-agnostic learning strategy trained exclusively on synthetic data generated from label maps (Hoffmann et al., 2021). TransMorph incorporated Vision Transformer architecture with a three-stage Swin Transformer encoder for hierarchical feature extraction (Chen et al., 2022). uniGradICON and uniGradICONiso represented foundation registration models using gradient inverse consistency regularization, with the latter including instance-specific optimization (Tian et al., 2024). VFA employed dual-stream encoders with vector field cross-attention decoders for pixel-level correspondence retrieval (Liu et al., 2024a). VoxelMorph served as the canonical U-Net-based learning framework for fast deformable registration (Balakrishnan et al., 2019). The top-performing methods in LUMIR employed diverse architectural strategies, ranging from specialized ConvNet designs to foundation model adaptations.
@@
Bailiang utilized a dual-stream encoder architecture with registration-specific components, employing coarse-to-fine optimization pyramids and progressive registration to handle large deformations through iterative composition of deformation fields. Correlation layers were used to establish voxel-wise correspondences between moving and fixed images (Jian et al., 2024).
@@
next-gen-nn implemented a robust backbone network (Zhang et al., 2024a) with several architectural enhancements, including co-attention mechanisms (Chen et al., 2021), large kernel convolutions, and increased convolutional channels in the deformation prediction stage. The method was further refined using image-guided bilateral filtering to improve registration accuracy.
@@
honkamj introduced a novel approach focusing on mathematical guarantees, implementing inverse consistency, symmetry, and topology preservation by construction rather than through loss terms. The method employed an auxiliary network formulation where $f(A,B)=u(A,B)\circ u(B,A)^{-1}$, ensuring cycle-consistency across a multi-resolution pipeline. Features were extracted separately before deformation computation, with symmetric halfway deformations composed at the highest resolution (Honkamaa and Marttinen, 2024).
@@
LoRA-FT addressed the computational challenges of foundation model adaptation by applying Low-rank Adaptation (LoRA (Hu et al., 2022)) to uniGradICON foundation models (Tian et al., 2024). This approach decomposed weight updates into low-rank matrices, reducing trainable parameters by orders of magnitude while enabling rapid domain adaptation without catastrophic forgetting or requiring anatomical label supervision.
@@
MadeForLife employed the GroupMorph (Tan et al., 2024) structure with dual encoders and specialized grouping update modules (GUMs). The method decomposed deformation fields into multiple subfields with different receptive fields, facilitated by grouping modules for feature correlation and contextual fusion modules for inter-group communication, enabling finer-grained optimization.
@@
lukasf refined a pre-trained TransMorph (Chen et al., 2022) model using the Fisher Adam (FAdam) optimizer (Hwang, 2024) for improved convergence stability and incorporated gradient correlation in the similarity measure to enhance deformation smoothness and anatomical alignment.
@@
TimH employed the ConstrICON framework (Greer et al., 2023) to achieve multistep inverse consistency through architectural design rather than penalty terms. The approach used three residual U-Nets in a recursive TwoStepConsistent (TSC) operator structure, parameterizing transformations via Lie groups. Stationary velocity fields were smoothed using average pooling, and selective Jacobian determinant regularization was applied to mitigate folding.
@@
VROC utilized a classical variational registration approach with task-specific preprocessing, including Gaussian smoothing and multi-Otsu thresholding to generate intensity-based label maps. The method performed initial deformable registration using normalized cross-correlation (NCC) forces on label maps, followed by Demons forces refinement on original intensity images for precise alignment (Werner et al., 2014; Ehrhardt et al., 2015).
@@
DutchMasters introduced a novel instance optimization algorithm addressing multi-objective optimization challenges in registration. The method applied gradient projection techniques to handle conflicting updates between similarity and regularization objectives, projecting conflicting gradients into common spaces when cosine similarity indicated gradient conflicts, while preserving non-conflicting gradient updates (Zhang et al., 2024b).

### v3

@@ Abstract
Medical image registration is critical for clinical applications, and fair benchmarking of different methods is essential for monitoring ongoing progress in the field. To date, the Learn2Reg 2020-2023 challenges have released several complementary datasets and established metrics for evaluations. Building on this foundation, the 2024 edition expands the challenge’s scope to cover a wider range of registration scenarios, particularly in terms of modality diversity and task complexity, by introducing three new tasks, including large-scale multi-modal registration and unsupervised inter-subject brain registration, as well as the first microscopy-focused benchmark within Learn2Reg. The new datasets also inspired new method developments, including invertibility constraints, pyramid features, keypoints alignment and instance optimisation. Visit Learn2Reg at https://learn2reg.grand-challenge.org.
@@ 1 Introduction
This paper serves as the official summary and documentation of the Learn2Reg 2024 challenge, held in conjunction with MICCAI 2024. As an overview paper, it introduces three new benchmark datasets addressing previously underrepresented registration scenarios, documents the methods developed by participating teams, and provides a comparative analysis of their performance across tasks. While this paper offers a comprehensive cross-task perspective on methods and results, individual task organizers may publish separate, in-depth analyses focusing on specific research questions, extended evaluations, or foundational model development. Our primary contributions in this challenge overview include: (1) establishing new publicly available benchmarks for multimodal intraoperative registration, large-scale unsupervised brain alignment, and histopathology microscopy registration; (2) documenting and comparing diverse methodological approaches across these tasks; and (3) identifying current limitations and future directions for medical image registration research through cross-task insights.
@@ 2.1 Medical Image Registration Challenges
Several notable medical image registration challenges have significantly advanced the field. The EMPIRE10 challenge focused on lung CT registration, offering comprehensive evaluations including landmark distances and Jacobian determinants (murphy2011evaluation). The CuRIOUS challenge addressed intra-operative ultrasound and MRI registration (xiao2019evaluation), while ANHIR specialized in histological image registration (borovec2020anhir), each attracting multiple participating teams and employing diverse evaluation metrics. The Continuous Registration Challenge, co-hosted with WBIR 2018, aimed to unify multiple registration tasks but was limited by its integration requirements, specifically the mandatory use of the SuperElastix C++ API framework (marstal2019continuous). More recently, the ACROBAT challenges in 2022 and 2023 addressed whole-slide image registration for breast cancer tissue, utilizing large datasets and providing valuable insights into algorithm performance (weitz2024acrobat). These challenges have collectively advanced the development and benchmarking of medical image registration algorithms, each contributing unique datasets and frameworks that have influenced Learn2Reg tasks and evaluation.
@@ 2.2 Learn2Reg to Date
Since its inception, the Learn2Reg challenge has served as a key platform for advancing medical image registration through open benchmarking and community collaboration. Over multiple editions, the challenge has brought together researchers worldwide to tackle diverse registration problems across various imaging modalities and anatomical sites.
The Learn2Reg series began with the Learn2Reg Tutorial at MICCAI 2019 in Shenzhen, which introduced the community to emerging deep learning approaches for medical image registration (learn2reg2019tutorial). This foundational event set the stage for the challenge’s collaborative and multidisciplinary focus.
The 2020 and 2021 editions of Learn2Reg were the initial official challenges, introducing six datasets covering multiple anatomies—including brain, abdomen, and thorax—and imaging modalities such as MRI, CT, and ultrasound. Both editions were held as community challenges with workshops at the respective MICCAI conferences, following a Type 1 challenge format with training images and labels provided, while test labels remained withheld for evaluation. These editions introduced the challenge’s open-source evaluation and ranking code, as well as baseline methods, enabling standardized and reproducible benchmarking. Together, they attracted over 20 international teams submitting more than 65 methods, offering comprehensive insights into the performance of conventional and deep-learning registration approaches across multiple clinically relevant tasks (hering2022learn2reg).
In 2022 and 2023, Learn2Reg held two further challenge workshops at MICCAI, focusing on continuation and expansion of the initial benchmarks. These editions introduced the NLST dataset (CT Lung Screening Registration Data) and pushed towards Type 3 challenges for selected tasks, where both training and test images, along with their labels, remain hidden during evaluation to better simulate real-world conditions.
In parallel, the OncoReg spin-off challenge was launched at MICCAI 2023, introducing the ThoraxCBCT dataset with longitudinal lung CBCT and FBCT scans from clinical settings. OncoReg targets oncological applications to advance AI-driven registration for monitoring lung cancer progression and improving radiotherapy planning (heyer2025oncoreg).
The 2024 edition continues this trajectory, expanding into new imaging domains including intraoperative multimodal imaging and microscopy, while addressing increasingly complex registration problems.
@@ 3.1 Challenge Organisation
The Learn2Reg 2024 challenge was organised by a dedicated core team led by Mattias Heinrich (University of Lübeck), who coordinated all aspects of the event, including website maintenance, evaluation infrastructure, and the associated workshop. The challenge consisted of three tasks, each developed and maintained by a separate expert team: ReMIND2Reg (Harvard Medical School, led by Reuben Dorent), LUMIR (Johns Hopkins University, led by Junyu Chen), and COMULISglobe SHG/BF (COMULISglobe community, led by Andreas Walter, Hochschule Aalen, (walter2021correlative; rudraiah2024correlated; walter2020correlated)). These tasks addressed diverse challenges in biomedical image registration and are described in detail in later sections.
Returning to a type 2 challenge format, Learn2Reg 2024 followed a two-phase structure hosted on the Grand Challenge platform [^1] https://learn2reg.grand-challenge.org/. During the validation phase, participants were given access to training and validation data to develop and test their methods locally. In the subsequent test phase, participants submitted their final algorithms, which were evaluated on hidden test datasets. The evaluation was performed using a standardized, open-source framework [^2] https://github.com/MDL-UzL/L2R, ensuring fairness, reproducibility, and transparency. Performance was assessed using task-specific quantitative metrics, with results published on public leaderboards. The ranking scheme is based on the ranking scheme of the Medical Decathlon (statistical comparisons of method performance across metrics, aggregated into overall task scores) (antonelli2022medical).
@@ 3.2.1 ReMIND2Reg
The challenge task is to register post-resection 3D iUS volumes to either contrast-enhanced T1-weighted (ceT1) or T2-weighted MRI, depending on availability for each case. Submitted methods must be able to handle missing modalities and account for large deformations, resections, and noise inherent in iUS data. The dataset is derived from the ReMIND collection (juvekar2023remind) at Brigham and Women’s Hospital. Specifically, the ReMIND2Reg dataset includes 99 training cases with 99 3D iUS, 93 ceT1, and 62 T2 scans, 4 validation cases with 4 3D US, 4 ceT1, and 4 T2, and 10 test cases with 10 3D US, 10 ceT1 and 10 T2.
@@
Full data and acquisition protocols are described in (juvekar2023remind). The dataset is shared under a CC BY 4.0 license and is available via TCIA [^3] https://doi.org/10.1101/2023.09.14.23295596.
@@ 3.2.2 LUMIR
Participants are required to submit deformation fields aligning specified image pairs. The performance is assessed via Dice similarity and Hausdorff-95 distance on segmentation labels, target registration error on anatomical landmarks, and the fraction of non-diffeomorphic volumes (NDV) (liu2024jacobian) to measure deformation plausibility. Baseline code and evaluation scripts are provided online [^4] https://github.com/JHU-MedImage-Reg/LUMIR_L2R. An introductory video is also available [^5] https://cloud.imi.uni-luebeck.de/s/tm9DEFiDa9XH35k/download/Learn2Reg_LUMIR_JChen.mp4.
This section documents LUMIR within the Learn2Reg 2024 challenge framework. As a multi-task overview, we focus on describing the task design, documenting participating methods, presenting comparative rankings, and identifying cross-task insights, particularly regarding the impact of dataset scale on classical versus deep learning approaches. The value of including LUMIR lies in enabling direct comparison with ReMIND2Reg and COMULISglobe, and providing a complete record of all Learn2Reg 2024 tasks. A complementary in-depth analysis with extensive ablation studies, zero-shot generalization experiments, and foundational model evaluation is provided in a dedicated paper (chen2025LUMIRchallengepathwayfoundational). Relevant datasets and prior methods are further detailed in (dufumier2022openbhb; taha2023magnetic; marcus2007oasis; liu2024jacobian).
@@ 3.2.3 COMULISglobe SHG/BF
Further details on the underlying imaging and dataset are available in (keikhosravi2020intensity). More information on the goals of the COMULISglobe initiative and the broader community can be found at the COMULISglobe website [^6] https://www.comulis.eu/.
@@ 3.2.4 NLST
The challenge consists of two phases. In Phase 1, participants developed and tested algorithms locally using 110 training and 22 validation cases, all sourced from the National Lung Screening Trial (NLST) (national2011reduced)  [^7] https://wiki.cancerimagingarchive.net/display/NLST/National+Lung+Screening+Trial. Each case consists of a pair of thoracic CT scans acquired either longitudinally (screening cohort) or across different respiratory phases (radiotherapy cohort). Automatic keypoint correspondences (heinrich2015estimating) and lung masks are provided to support model development. In Phase 2, the top-performing teams were invited to submit training Docker containers for remote execution on a hidden, extended dataset comprising 1300 training, over 100 validation, and more than 200 test cases. This second part of the dataset includes additional annotations not publicly released.
@@ 4.1.1 ReMIND2Reg
Figure 3: Results for the ReMIND2Reg task. Target Registration Error (TRE) on manual landmarks and TRE30 (the TRE computed on the 30% of landmarks with the largest initial errors) are reported for all methods. * Baseline method provided by organizers.
@@
During the validation phase of the ReMIND2Reg task, a total of thirteen teams submitted results, along with two baseline methods provided by the challenge organizers. The task’s complexity was reflected in the exceptionally high number of validation entries, with 253 submissions indicating that teams required extensive iterative optimization to handle the challenging multimodal ultrasound-to-MRI registration problem. In the test phase, participation reduced to five teams plus one baseline method. This reduction highlights the technical difficulty of registering post-resection intraoperative ultrasound to preoperative MRI, particularly given the presence of brain shift and tissue deformation. As a baseline, niftyreg was provided for the ReMIND2Reg test phase. The submitting teams included junyi-wang, VROC, next-gen-nn, IGRTS, and SamPLe.
@@
niftyreg, which employed the NiftyReg open-source software package [^8] https://github.com/KCL-BMEIS/niftyreg/wiki implementing a symmetric block-matching registration approach. This method used iterative block-matching to establish point correspondences between ultrasound and MRI images, followed by least trimmed squares (LTS) regression to determine transformation parameters. To identify informative regions, the algorithm divided images into uniform blocks and selected those with the highest intensity variance. Block similarity was assessed using normalized cross-correlation, and matching was performed based on these scores. To ensure inverse consistency, the method averaged directional transformation matrices at each iteration, maintaining symmetry throughout the registration process (drobny2018registration).
@@
junyi-wang developed a coarse-to-fine pipeline combining imaging style transfer with traditional registration techniques. Their approach used 3D CycleGAN to generate synthetic T1-style images from ultrasound data, enabling unified signal distributions across modalities. The registration process employed hierarchical block matching via NiftyReg for affine transformation estimation, followed by a pretrained SynthMorph model for local deformation inference around resection areas. VROC employed its fully differentiable variational registration framework (werner2014estimation; ehrhardt2015variational) [^9] https://github.com/IPMI-ICNS-UKE/vroc, combining task-specific preprocessing with rigid registration optimization. For the ReMIND2Reg task, VROC applied Gaussian smoothing and threshold-based clipping of ultrasound intensity values during preprocessing. The registration pipeline performed initial vector field estimation via global normalized cross-correlation (NCC) loss optimization, followed by fine-tuning using normalized gradient fields (NGF) loss optimization to handle the challenging multimodal ultrasound-to-MRI registration scenario.
@@
next-gen-nn addressed the ReMIND2Reg challenge using an unsupervised multimodal registration method based on Multilevel Correlation Balanced Optimization (MCBO), building upon the ConvexAdam framework (siebert2024convexadam). The approach was specifically designed to handle two key difficulties: the absence of supervised labels and the presence of large deformations between pre-operative MRI and intra-operative ultrasound images. For paired multimodal moving images (ceT1 and T2), the method adopted maximum fusion strategies to optimize registration performance across the challenging modality gap. IGRTS developed a topological higher-order Markov Random Field approach for deformable registration. Their method utilized multiscale multimodal Gabor attribute vectors to characterize MRI and ultrasound images, with the 3D deformation field’s topological framework constructed through Jacobian determinants derived from higher-order cliques. The optimization employed a multi-resolution QPBO (Quadratic Pseudo-Boolean Optimization) strategy to handle the complex higher-order MRF energy function.
@@ 4.1.2 LUMIR
Figure 4: Results for the LUMIR task. The table presents final test set results for all participating teams and baseline methods. Reported metrics include Target Registration Error (TRE) on manual landmarks, Dice-Sørensen coefficient (DSC) and 95th percentile Hausdorff distance (HD95) on brain label maps, as well as the non-diffeomorphic volume (NDV) of displacement fields. The accompanying box plot provides a visual overview of DSC performance distribution across all methods. * Baseline method provided by organizers.
@@
A distinctive aspect of the LUMIR challenge was its extensive post-challenge analysis, documented in detail in ”Beyond the LUMIR challenge: The pathway to foundational registration models” (chen2025LUMIRchallengepathwayfoundational). This companion paper provides comprehensive validation with particular emphasis on foundational registration models, zero-shot evaluation, and out-of-domain generalization tasks. The present section summarizes the key findings and participating methods, while readers are referred to the companion paper for in-depth technical analysis and broader implications for developing generalizable brain MRI registration frameworks. The LUMIR challenge included eight baseline methods spanning classical optimization-based approaches and modern deep learning frameworks. ANTsSyN employed symmetric, time-varying diffeomorphic registration using cross-correlation within the space of diffeomorphic maps (avants2008symmetric). deedsBCV utilized discrete optimization with Markov random field strategies and minimum spanning trees to handle sliding motions efficiently (heinrich2012mind). SynthMorph implemented a contrast-agnostic learning strategy trained exclusively on synthetic data generated from label maps (hoffmann2021synthmorph). TransMorph incorporated Vision Transformer architecture with a three-stage Swin Transformer encoder for hierarchical feature extraction (chen2022transmorph). uniGradICON and uniGradICONiso represented foundation registration models using gradient inverse consistency regularization, with the latter including instance-specific optimization (tian2024unigradicon). VFA employed dual-stream encoders with vector field cross-attention decoders for pixel-level correspondence retrieval (liu2024vector). VoxelMorph served as the canonical U-Net-based learning framework for fast deformable registration (balakrishnan2019voxelmorph). The top-performing methods in LUMIR employed diverse architectural strategies, ranging from specialized ConvNet designs to foundation model adaptations.
@@
Bailiang utilized a dual-stream encoder architecture with registration-specific components, employing coarse-to-fine optimization pyramids and progressive registration to handle large deformations through iterative composition of deformation fields. Correlation layers were used to establish voxel-wise correspondences between moving and fixed images (jian2024mamba).
@@
next-gen-nn implemented a robust backbone network (zhang2024memwarp) with several architectural enhancements, including co-attention mechanisms (chen2021car), large kernel convolutions, and increased convolutional channels in the deformation prediction stage. The method was further refined using image-guided bilateral filtering to improve registration accuracy.
@@
honkamj introduced a novel approach focusing on mathematical guarantees, implementing inverse consistency, symmetry, and topology preservation by construction rather than through loss terms. The method employed an auxiliary network formulation where $f(A,B)=u(A,B)\circ u(B,A)^{-1}$, ensuring cycle-consistency across a multi-resolution pipeline. Features were extracted separately before deformation computation, with symmetric halfway deformations composed at the highest resolution (honkamaa2023sitreg).
@@
LoRA-FT addressed the computational challenges of foundation model adaptation by applying Low-rank Adaptation (LoRA (hu2022lora)) to uniGradICON foundation models (tian2024unigradicon). This approach decomposed weight updates into low-rank matrices, reducing trainable parameters by orders of magnitude while enabling rapid domain adaptation without catastrophic forgetting or requiring anatomical label supervision.
@@
MadeForLife employed the GroupMorph (tan2024groupmorph) structure with dual encoders and specialized grouping update modules (GUMs). The method decomposed deformation fields into multiple subfields with different receptive fields, facilitated by grouping modules for feature correlation and contextual fusion modules for inter-group communication, enabling finer-grained optimization.
@@
lukasf refined a pre-trained TransMorph (chen2022transmorph) model using the Fisher Adam (FAdam) optimizer (hwang2024fadam) for improved convergence stability and incorporated gradient correlation in the similarity measure to enhance deformation smoothness and anatomical alignment.
@@
TimH employed the ConstrICON framework (greer2023inverse) to achieve multistep inverse consistency through architectural design rather than penalty terms. The approach used three residual U-Nets in a recursive TwoStepConsistent (TSC) operator structure, parameterizing transformations via Lie groups. Stationary velocity fields were smoothed using average pooling, and selective Jacobian determinant regularization was applied to mitigate folding.
@@
VROC utilized a classical variational registration approach with task-specific preprocessing, including Gaussian smoothing and multi-Otsu thresholding to generate intensity-based label maps. The method performed initial deformable registration using normalized cross-correlation (NCC) forces on label maps, followed by Demons forces refinement on original intensity images for precise alignment (werner2014estimation; ehrhardt2015variational).
@@
DutchMasters introduced a novel instance optimization algorithm addressing multi-objective optimization challenges in registration. The method applied gradient projection techniques to handle conflicting updates between similarity and regularization objectives, projecting conflicting gradients into common spaces when cosine similarity indicated gradient conflicts, while preserving non-conflicting gradient updates (zhang2024improving).